OpenStudy (anonymous):

Write a formula for this sequence: 1/5, 3/5, 1, 1 2/5... A. -1/5 + 2/5 n B. 1/5 + 2/5 n C. 1 2/5 - 2/5 n

OpenStudy (anonymous):

how do you do it

OpenStudy (anonymous):

A if n=1:(-1/5)+(2/5)=1/5 n=2:(-1/5)+(4/5)=3/5 n=3:(-1/5)+(6/5)=(5/5)=1

OpenStudy (anonymous):

umm i am lost

OpenStudy (anonymous):

you put values for n in the general formula,here -1/5 + 2/5 n for each value of n you get a term of the sequence 1/5, 3/5, 1, 1 2/5...

OpenStudy (anonymous):

o ok i get it know. Thank you